# CHIP: Adaptive Compliance for Humanoid Control through Hindsight Perturbation

### v1

@@ I Introduction
RL-based compliance controllers have shown promising results on quadruped robots [unifp, facet] by tracking motion spring-damper-like reference motion. However, these methods require substantial effort to generate large volumes of synthetic data that emulate spring-damper end-effector dynamics. Extending these approaches to humanoids is challenging, as it is difficult to ensure that such synthetic perturbation data remains within the distribution of natural human motion. In parallel, advances in RL-based motion tracking [beyondmimic, twist2, any2track] have demonstrated that learning from human demonstrations is a scalable approach for constructing general-purpose humanoid controllers. However, these methods do not provide a principled way to integrate compliant or impedance control into humanoid tracking frameworks. In this work, we pose the central question: Can we develop a scalable and generalizable approach for reconciling the high-gain stiffness required for humanoid agile motions with the variable compliance needed for safe contact-rich manipulation? Some concurrent efforts [gentlehumanoid, softmimic] attempt to extend FACET [facet] to humanoids by augmenting retargeted human motion data with synthetic observations per responses under force perturbations, and then relying on tracking rewards to encourage imitation of the augmented trajectories. Yet, as is common in RL-based motion tracking [hover, omnih2o, twist2, amo, head], reward computation typically depends on rich reference information such as link poses, velocities, joint positions, and joint velocities, which are challenging to modify reliably during data augmentation.
@@ II-A Variable and adaptive compliance manipulation
Adaptive compliance control has been used in manipulation to handle contact uncertainty and solve tasks that require force control. Prior work [vices, acp] shows that variable impedance control could enable a robot to perform tasks that require significant contact force control, such as door opening, or to maintain surface contact, such as wiping. Under contact uncertainty, an adaptive compliance controller can also help robots perform stable grasps [springgrasp, hmc]. Most prior adaptive compliance control studies use a tabletop arm with a model-based impedance controller. In our work, we draw inspiration from task design in these works, but apply it to a learned adaptive compliance controller on a humanoid robot.
@@ II-B Compliant control of legged robot
Recent success in legged locomotion and loco-manipulation benefits from large-scale reinforcement learning in simulations. RL motion tracking has become a pivotal task enabling general humanoid control. Policies trained purely for motion tracking, often inspired by frameworks such as DeepMimic [deepmimic], tend to be stiff and fragile under force perturbations. When the objective treats any deviation from a reference motion as an error that must be aggressively corrected, the robot generates large, uncontrolled forces during unexpected contact. To address this, recent research has pursued two distinct paradigms: learning to resist external forces or learning to comply with them.
@@
Learning to Resist Disturbances. One line of work develops controllers designed to maintain tracking performance despite large, unknown external forces. These methods typically train policies with randomized external pushes to encourage robustness. Approaches that combine motion-tracking rewards with random external forces instruct the robot to adhere to the reference trajectory despite interaction forces [ze2025twist, fu2022dwbc, zhang2025falcon]. A prominent example is FALCON [zhang2025falcon], which uses a torque-aware 3D force curriculum, enabling the robot to gradually learn in simulation to resist strong disturbances and pull heavy carts. While effective for applications requiring forceful opposition, this stiffness-centric approach is ill-suited for contact-rich manipulation scenarios—such as robot-human collaboration, wiping, or handling fragile objects—which require the robot to yield in a controlled, spring-like manner.
@@
Learning to Comply with Interactions. In contrast to force rejection, a second line of research aims to learn controllers that yield to external forces. One approach creates impedance behaviors ”from scratch” using RL. For example, UniFP [unifp] and FACET [facet] train unified policies to mimic target spring-mass-damper dynamics. UniFP learns a force estimator for admittance control, while FACET modulates policy stiffness as a control input, enabling tasks such as wiping using only proprioception. However, because they are primarily designed for quadruped robots with a single end-effector, scaling these methods to humanoids is challenging because generating synthetic compliance data that matches the distribution of natural human movement is difficult.
@@
To sidestep the difficulty of discovering behaviors from scratch, prior work has attempted to build compliance into motion-tracking frameworks. SoftMimic [softmimic] proposes a learning-from-examples framework, using an offline inverse kinematics (IK) solver to generate an augmented dataset of stylistically compliant motions. An RL policy then learns to reproduce these pre-authored responses. Similarly, GentleHumanoid [gentlehumanoid] augments keypoint reference motions with desired spring-damper dynamics during reward calculation. However, this creates an optimization conflict between motion tracking rewards and compliance objectives, often leading to degraded tracking accuracy or reduced agility. To make things worse, relying on offline data augmentation or reward tuning limits their applicability for scaling up to large, diverse motion datasets.
@@
Our work builds upon these insights but proposes a more direct and scalable alternative. We introduce CHIP, a training recipe that integrates principled impedance-control objectives directly into the online RL loop via hindsight perturbation, without compromising tracking performance. CHIP obviates the need for the extensive synthetic interaction curriculum [unifp, facet] or the offline motion augmentation pipelines of SoftMimic [softmimic] and GentleHumanoid [gentlehumanoid]. By specifically targeting modifications in the input space, our method offers a simple yet powerful way to equip standard humanoid tracking systems with adaptive compliance, narrowing the gap between agile motion and safe, controllable physical interaction.
@@ II-C Humanoid control interfaces
Recent advances in reinforcement learning for humanoid whole-body control have produced remarkable results in agile locomotion [beyondmimic, kungfubot, hub, sonic] and whole-body manipulation [ze2025twist, amo, homie]. Extending motion-tracking approaches, several methods [ze2025twist, gmt] use human whole-body motion—encompassing both link- and joint-level states—to drive humanoid robot control. However, such control interfaces are complicated to use for both teleoperation and specifying high-level task commands. To simplify humanoid control, several methods [homie, amo] adopt a decoupled upper-lower-body policy, in which the robot receives root-velocity and height commands for locomotion and upper-body target-motion commands for manipulation. Other approaches explore keypoint-based control [omnih2o, hover, sonic, head, twist2, clone], enabling intuitive teleoperation by mapping head and wrist movements from a VR headset to humanoid motion. CHIP upgrades humanoid keypoint control with an adaptive-compliance module. It is compatible with both local keypoint tracking, as in SONIC [sonic], OmniH2O [omnih2o], and global keypoint tracking [clone, head], while additionally providing a plug-and-play module that directly integrates task-aware control into existing keypoint-based frameworks.
@@ III Method
Figure 2: Overview of training and deployment of CHIP: this general formulation takes tracking targets, end-effector compliance coefficients, and state history as default inputs. For the local tracking policy, we provide an additional lower-body reference pose, generated via the kinematic planner used in SONIC [sonic]. It adapts end-effector compliance based on the input compliance coefficient and force implicitly estimated from proprioception history. During training, the policy observes recovered keypoint poses by removing the hindsight perturbation, and tracking rewards are calculated from the original reference motion. The trained adaptive compliance policy could be controlled via multiple interfaces, such as VR teleoperation, a kinematic planner, and a vision-language-action (VLA) model.
Figure 3: Different formulations for adaptive compliant tracking policy training. (a) Ours (CHIP), the policy observes hindsight tracking goal, a modified end-effectors goal in the opposite direction of the perturbation force. (b) Modify keypoint goals for reward as in GentleHumanoid [gentlehumanoid]. (c) Edit entire kinematic reference motion to account for end-effector perturbation as in SoftMimic [softmimic].
@@
Prior work [unifp, facet] and recent concurrent work [gentlehumanoid, softmimic] train policies $\pi(\bm{s},\bm{g})$ that observe tracking goals $\bm{g}$ from the original reference motion, but use tracking goals modified by perturbations, $\bm{g}_{\text{ptb}}$, to calculate the reward that encourages the robot to track the desired compliant response under force perturbations rather than the original motions (Fig. 3(b)):
@@
This formulation streamlines training while enhancing robustness and agility. Unlike methods such as [unifp, facet], which explicitly learn an external force estimator, our approach uses a single policy model optimized solely with the PPO objective. The policy implicitly acquires force awareness through proprioceptive observations and past actions. To further improve sensitivity to external perturbations during training, we provide the critic with the ground-truth external force as a privileged observation. In addition, we supply both the actor and critic with a 10-step history of proprioception and past actions, allowing the policy to infer perturbation-related information from noisy observations. At test time, the policy only receives the tracking goal $\bm{g}$ and proprioception. It implicitly estimates the perturbation force and generates actions that counteract disturbances according to the commanded compliance. To enable compliant following, similar to FACET [facet], we also implement a damper model on the target end-effector pose. Under an applied force, the end-effector pose $\bm{x}_{\text{eef}}$ deviates from the previous target $\bm{g}_{t-1}$; we update the target using:
@@ IV-A Compliant multi-robot grasping
To obtain a two-robot grasp, we extend the dexterous compliant grasp generation algorithm SpringGrasp [springgrasp] to a multi-humanoid collaborative grasp setting. We solve an optimization problem to obtain the 6D head poses and 3D wrist positions of two robots:
@@
After the object has been lifted, we can simultaneously translate the 3-point target poses of both robots to move the grasped large object. For multi-robot collaboration to work in diverse environments, we require high-frequency, infrastructure-free robot localization. Following BeyondMimic [beyondmimic], we combine:
@@
  • Fast-LIO2 based pose updates at 10 Hz [fastlio2],
@@ IV-B1 Local 3-point tracking policy
For tasks that do not require global localization, the compliant local 3-point policy is straightforward to deploy. Following SONIC [sonic], the robot is teleoperated using VR-tracked wrist poses along with root velocity and height commands from VR joysticks as shown in Fig. 4. Lower-body control is handled by a kinematic planner driven by these commands. The kinematic planner takes root velocity, angular velocity, and height as commands and outputs planned joint positions and velocities. Compliance coefficients for each end-effector are adjusted by the operator in VR. Teleoperation sessions collect paired action data and compliance coefficients from the teleoperation interface, as well as binocular RGB observations from the robot’s egocentric OAK camera. These data are used to finetune the GR00T N1.5 [groot] VLA model.
@@ V Experiments and Results
We compared CHIP with two baselines: 1) FALCON [zhang2025falcon], a policy that does not observe compliance but was trained with force perturbation, and 2) No force, a standard tracking policy that neither observes compliance nor was trained with force perturbation.
@@ V-A Experiment setup
We used Unitree G1 as our testing platform, and an OAK-D W camera was mounted on the robot head to provide egocentric binocular vision. All policies were trained on 64 Nvidia L40S GPUs, with 4096 environments per GPU, for 4 days. During deployment, the local 3-point tracking policy, kinematic motion planner, and teleoperation server were running on Unitree G1 Jetson NX onboard computer with TensorRT acceleration. The global 3-point tracking policy, together with the state estimator, was running on a desktop machine with an i7-13700K CPU and an RTX 3090 GPU. We use the same desktop computer to run VLA policies.
@@ V-B Tracking performance of CHIP
Ours ($\frac{1}{k}$ = 0.0) 0.08 (0.04) 0.09 (0.04) 0.02 (0.02) 0.09 (0.06)
Ours ($\frac{1}{k}$ = 0.02) 0.08 (0.05) 0.08 (0.05) 0.02 (0.02) 0.10 (0.06)
Ours ($\frac{1}{k}$ = 0.05) 0.08 (0.04) 0.08 (0.04) 0.02 (0.02) 0.11 (0.06)
@@
TABLE I: Tracking errors are measured in the absence of external forces for our method under three distinct evaluation-time settings, and compared against two baseline approaches. Entries are reported as mean (standard deviation). Metrics: $\epsilon_{p}^{g}$ – global position error [m], $\epsilon_{r}^{g}$ – global orientation error [rad], $\epsilon_{p}^{l}$ – local position error [m], and $\epsilon_{r}^{l}$ – local orientation error [rad].
We analyze the tracking accuracy and the robot’s response to force perturbations under different end-effector compliance settings. Fig. I shows the tracking accuracy measured over 100 trajectories sampled from the TWIST dataset [ze2025twist]; it shows that the use of hindsight perturbation training yields position tracking on par with the no-compliance and no-force-perturbation baseline, and only slightly affects orientation tracking performance. Meanwhile, the results show that varying end-effector stiffness does not significantly affect tracking performance, indicating that the policy can distinguish between the joint’s internal driving force required for agile performance and the external perturbation force. We also measured how the compliance coefficient affects the robot’s responses to force perturbations. Fig. 6 shows that for CHIP, displacement increases linearly while increasing the compliance coefficient. For the baseline, which naively adds a reward for tracking the modified end-effector trajectory, the robot is less responsive to changes in the compliance coefficient because of conflicts among the tracking reward terms.
Figure 6: EEF displacement under 20 N external force in global tracking setting. Reward Tuning: baseline that adds a modified end-effector tracking reward but removes the hindsight perturbation mechanism.

### v2

@@ I Introduction
RL-based compliance controllers have shown promising results on quadruped robots [27, 20] by tracking motion spring-damper-like reference motion. However, these methods require substantial effort to generate large volumes of synthetic data that emulate spring-damper end-effector dynamics. Extending these approaches to humanoids is challenging, as it is difficult to ensure that such synthetic perturbation data remains within the distribution of natural human motion. In parallel, advances in RL-based motion tracking [12, 23, 26] have demonstrated that learning from human demonstrations is a scalable approach for constructing general-purpose humanoid controllers. However, these methods do not provide a principled way to integrate compliant or impedance control into humanoid tracking frameworks. In this work, we pose the central question: Can we develop a scalable and generalizable approach for reconciling the high-gain stiffness required for humanoid agile motions with the variable compliance needed for safe contact-rich manipulation? Some concurrent efforts [13, 15] attempt to extend FACET [20] to humanoids by augmenting retargeted human motion data with synthetic observations per responses under force perturbations, and then relying on tracking rewards to encourage imitation of the augmented trajectories. Yet, as is common in RL-based motion tracking [8, 7, 23, 10, 4], reward computation typically depends on rich reference information such as link poses, velocities, joint positions, and joint velocities, which are challenging to modify reliably during data augmentation.
@@ II-A Variable and adaptive compliance manipulation
Adaptive compliance control has been used in manipulation to handle contact uncertainty and solve tasks that require force control. Prior work [16, 9] shows that variable impedance control could enable a robot to perform tasks that require significant contact force control, such as door opening, or to maintain surface contact, such as wiping. Under contact uncertainty, an adaptive compliance controller can also help robots perform stable grasps [3, 18]. Most prior adaptive compliance control studies use a tabletop arm with a model-based impedance controller. In our work, we draw inspiration from task design in these works, but apply it to a learned adaptive compliance controller on a humanoid robot.
@@ II-B Compliant control of legged robot
Recent success in legged locomotion and loco-manipulation benefits from large-scale reinforcement learning in simulations. RL motion tracking has become a pivotal task enabling general humanoid control. Policies trained purely for motion tracking, often inspired by frameworks such as DeepMimic [17], tend to be stiff and fragile under force perturbations. When the objective treats any deviation from a reference motion as an error that must be aggressively corrected, the robot generates large, uncontrolled forces during unexpected contact. To address this, recent research has pursued two distinct paradigms: learning to resist external forces or learning to comply with them.
@@
Learning to Resist Disturbances. One line of work develops controllers designed to maintain tracking performance despite large, unknown external forces. These methods typically train policies with randomized external pushes to encourage robustness. Approaches that combine motion-tracking rewards with random external forces instruct the robot to adhere to the reference trajectory despite interaction forces [22, 6, 25]. A prominent example is FALCON [25], which uses a torque-aware 3D force curriculum, enabling the robot to gradually learn in simulation to resist strong disturbances and pull heavy carts. While effective for applications requiring forceful opposition, this stiffness-centric approach is ill-suited for contact-rich manipulation scenarios—such as robot-human collaboration, wiping, or handling fragile objects—which require the robot to yield in a controlled, spring-like manner.
@@
Learning to Comply with Interactions. In contrast to force rejection, a second line of research aims to learn controllers that yield to external forces. One approach creates impedance behaviors ”from scratch” using RL. For example, UniFP [27] and FACET [20] train unified policies to mimic target spring-mass-damper dynamics. UniFP learns a force estimator for admittance control, while FACET modulates policy stiffness as a control input, enabling tasks such as wiping using only proprioception. However, because they are primarily designed for quadruped robots with a single end-effector, scaling these methods to humanoids is challenging because generating synthetic compliance data that matches the distribution of natural human movement is difficult.
@@
To sidestep the difficulty of discovering behaviors from scratch, prior work has attempted to build compliance into motion-tracking frameworks. SoftMimic [15] proposes a learning-from-examples framework, using an offline inverse kinematics (IK) solver to generate an augmented dataset of stylistically compliant motions. An RL policy then learns to reproduce these pre-authored responses. Similarly, GentleHumanoid [13] augments keypoint reference motions with desired spring-damper dynamics during reward calculation. However, this creates an optimization conflict between motion tracking rewards and compliance objectives, often leading to degraded tracking accuracy or reduced agility. To make things worse, relying on offline data augmentation or reward tuning limits their applicability for scaling up to large, diverse motion datasets.
@@
Our work builds upon these insights but proposes a more direct and scalable alternative. We introduce CHIP, a training recipe that integrates principled impedance-control objectives directly into the online RL loop via hindsight perturbation, without compromising tracking performance. CHIP obviates the need for the extensive synthetic interaction curriculum [27, 20] or the offline motion augmentation pipelines of SoftMimic [15] and GentleHumanoid [13]. By specifically targeting modifications in the input space, our method offers a simple yet powerful way to equip standard humanoid tracking systems with adaptive compliance, narrowing the gap between agile motion and safe, controllable physical interaction.
@@ II-C Humanoid control interfaces
Recent advances in reinforcement learning for humanoid whole-body control have produced remarkable results in agile locomotion [12, 19, 24, 14] and whole-body manipulation [22, 10, 1]. Extending motion-tracking approaches, several methods [22, 5] use human whole-body motion—encompassing both link- and joint-level states—to drive humanoid robot control. However, such control interfaces are complicated to use for both teleoperation and specifying high-level task commands. To simplify humanoid control, several methods [1, 10] adopt a decoupled upper-lower-body policy, in which the robot receives root-velocity and height commands for locomotion and upper-body target-motion commands for manipulation. Other approaches explore keypoint-based control [7, 8, 14, 4, 23, 11], enabling intuitive teleoperation by mapping head and wrist movements from a VR headset to humanoid motion. CHIP upgrades humanoid keypoint control with an adaptive-compliance module. It is compatible with both local keypoint tracking, as in SONIC [14], OmniH2O [7], and global keypoint tracking [11, 4], while additionally providing a plug-and-play module that directly integrates task-aware control into existing keypoint-based frameworks.
@@ III Method
Figure 2: Overview of training and deployment of CHIP: this general formulation takes tracking targets, end-effector compliance coefficients, and state history as default inputs. For the local tracking policy, we provide an additional lower-body reference pose, generated via the kinematic planner used in SONIC [14]. It adapts end-effector compliance based on the input compliance coefficient and force implicitly estimated from proprioception history. During training, the policy observes recovered keypoint poses by removing the hindsight perturbation, and tracking rewards are calculated from the original reference motion. The trained adaptive compliance policy could be controlled via multiple interfaces, such as VR teleoperation, a kinematic planner, and a vision-language-action (VLA) model.
Figure 3: Different formulations for adaptive compliant tracking policy training. (a) Ours (CHIP), the policy observes hindsight tracking goal, a modified end-effectors goal in the opposite direction of the perturbation force. (b) Modify keypoint goals for reward as in GentleHumanoid [13]. (c) Edit entire kinematic reference motion to account for end-effector perturbation as in SoftMimic [15].
@@
Prior work [27, 20] and recent concurrent work [13, 15] train policies $\pi(\bm{s},\bm{g})$ that observe tracking goals $\bm{g}$ from the original reference motion, but use tracking goals modified by perturbations, $\bm{g}_{\text{ptb}}$, to calculate the reward that encourages the robot to track the desired compliant response under force perturbations rather than the original motions (Fig. 3(b)):
@@
This formulation streamlines training while enhancing robustness and agility. Unlike methods such as [27, 20], which explicitly learn an external force estimator, our approach uses a single policy model optimized solely with the PPO objective. The policy implicitly acquires force awareness through proprioceptive observations and past actions. To further improve sensitivity to external perturbations during training, we provide the critic with the ground-truth external force as a privileged observation. In addition, we supply both the actor and critic with a 10-step history of proprioception and past actions, allowing the policy to infer perturbation-related information from noisy observations. At test time, the policy only receives the tracking goal $\bm{g}$ and proprioception. It implicitly estimates the perturbation force and generates actions that counteract disturbances according to the commanded compliance. To enable compliant following, similar to FACET [20], we also implement a damper model on the target end-effector pose. Under an applied force, the end-effector pose $\bm{x}_{\text{eef}}$ deviates from the previous target $\bm{g}_{t-1}$; we update the target using:
@@ IV-A Compliant multi-robot grasping
To obtain a two-robot grasp, we extend the dexterous compliant grasp generation algorithm SpringGrasp [3] to a multi-humanoid collaborative grasp setting. We solve an optimization problem to obtain the 6D head poses and 3D wrist positions of two robots:
@@
After the object has been lifted, we can simultaneously translate the 3-point target poses of both robots to move the grasped large object. For multi-robot collaboration to work in diverse environments, we require high-frequency, infrastructure-free robot localization. Following BeyondMimic [12], we combine:
@@
  • Fast-LIO2 based pose updates at 10 Hz [21],
@@ IV-B1 Local 3-point tracking policy
For tasks that do not require global localization, the compliant local 3-point policy is straightforward to deploy. Following SONIC [14], the robot is teleoperated using VR-tracked wrist poses along with root velocity and height commands from VR joysticks as shown in Fig. 7. Lower-body control is handled by a kinematic planner driven by these commands. The kinematic planner takes root velocity, angular velocity, and height as commands and outputs planned joint positions and velocities. Compliance coefficients for each end-effector are adjusted by the operator in VR. Teleoperation sessions collect paired action data and compliance coefficients from the teleoperation interface, as well as binocular RGB observations from the robot’s egocentric OAK camera. These data are used to finetune the GR00T N1.5 [2] VLA model.
@@ V Experiments and Results
  • Q5: What is the advantage of the continuous compliance command space used in CHIP?
@@ V-A Experiment setup
We used the Unitree G1 as our test platform, and an OAK-D W camera was mounted on the robot’s head to provide egocentric binocular vision. All policies were trained for 4 days on 64 Nvidia L40S GPUs, with 4096 environments per GPU. During deployment, the local 3-point tracking policy, kinematic motion planner, and teleoperation server were running on the Unitree G1 Jetson NX onboard computer with TensorRT acceleration. The global 3-point tracking policy, together with the state estimator, was running on a desktop machine with an i7-13700K CPU and an RTX 3090 GPU. We use the same desktop computer to run VLA policies.
@@ V-B Tracking performance of CHIP
Ours ($1/k=0.00$) 0.08 (0.04) 0.09 (0.04) 0.02 (0.02) 0.09 (0.06)
Ours ($1/k=0.02$) 0.08 (0.05) 0.08 (0.05) 0.02 (0.02) 0.10 (0.06)
Ours ($1/k=0.05$) 0.08 (0.04) 0.08 (0.04) 0.02 (0.02) 0.11 (0.06)
Gentlehumanoid [stiff: max force=15N] - - 0.04 (0.03) 0.15 (0.08)
Gentlehumanoid [compliant: max force=5N] - - 0.05 (0.04) 0.16 (0.10)
@@
TABLE I: Tracking errors are measured in the MuJoCo simulator without external force perturbation. The No perturb force baseline does not train with force perturbation at all; the FALCON [25] is trained with force perturbations on the end-effector but does not observe compliance. Gentlehumanoid [13] observes compliance commands but uses reward-shaping to get the compliant behavior. Entries are reported as mean (standard deviation). Metrics: For global policy $\epsilon_{p}^{g}$ – global position error [m], $\epsilon_{r}^{g}$ – global orientation error [rad]; for local policy $\epsilon_{p}^{l}$ – local position error [m], and $\epsilon_{r}^{l}$ – local orientation error [rad].
We analyze the tracking accuracy and the robot’s response to force perturbations under different end-effector compliance settings. Tab. I shows the tracking accuracy measured over 100 trajectories sampled from the TWIST dataset [22]; it shows that the use of hindsight perturbation training yields position tracking on par with the no-compliance (FALCON [25]) and no-force-perturbation baseline, with only a slight impact on orientation tracking performance. Meanwhile, Tab. I shows that varying end-effector stiffness does not significantly affect the tracking performance of our method, indicating that the policy can distinguish between the internal joint’s driving force required for agile performance and external perturbation forces. Compared with another compliant baseline, Gentlehumanoid [13], which removes dense joint-level tracking reward to achieve compliance behavior, CHIP outperforms it in terms of higher tracking accuracy. This result clearly shows that the design choice in CHIP, which modifies observations via hindsight perturbation rather than reward tuning, is a better approach to achieve both compliance behavior and accurate general tracking performance.